# Digital FAST: An AI-Driven Multimodal Framework for Rapid and Early Stroke Screening

### v1

@@
Abstract
@@
I Introduction
Stroke refers to a group of cerebrovascular disorders caused by the formation of blood clots or hemorrhagic events that obstruct or rupture blood vessels supplying the brain. Among these conditions, acute ischemic stroke is one of the leading causes of long-term disability and mortality worldwide[10, 12], with epidemiological studies estimating that approximately one in four individuals will experience a stroke during their lifetime[7]. Consequently, timely and appropriate intervention is of critical importance. Early detection of stroke is essential to enable prompt preventive measures and effective treatment. Notably, when individuals exhibit early signs of stroke, bystanders often lack sufficient medical knowledge to recognize these symptoms as accurately as healthcare professionals. This gap highlights the urgent need for automated methods capable of detecting stroke risk in real-world settings, particularly for non-expert users. In the context of rapid digitalization, artificial intelligence (AI) has significantly impacted the healthcare domain. Deep learning models constitute the foundation of various applications, including automated healthcare systems, clinical decision support, and medical diagnosis, where AI has, in certain cases, achieved performance superior to that of human specialists[1]. Multimodal learning approaches, which integrate information from multiple data sources, have demonstrated superior performance compared to models relying on a single modality[3]. However, in light of the rapid digitalization of healthcare and the growing adoption of multimodal deep learning techniques, there remains a notable gap between methodological advances and their application to practical, high-impact clinical scenarios such as prehospital stroke screening. Although prior studies have demonstrated the advantages of integrating heterogeneous data sources, multimodal stroke detection particularly under subject independent and prehospital settings has not yet been sufficiently explored[15]. Addressing this gap requires not only effective multimodal modeling strategies but also datasets and experimental settings that reflect real-world constraints.
@@ I Introduction
The remainder of this paper is organized as follows. Section II reviews related work on stroke detection and multimodal deep learning approaches. Section III presents the proposed multimodal framework, including the modality specific encoders and the fusion strategy. Section IV describes the dataset and data acquisition process in detail. Section V reports the experimental setup and quantitative performance results. Section VI provides ablation studies to analyze the contribution of individual modalities and fusion strategies. Section VII discusses the experimental findings, limitations, and practical implications of the proposed method. Finally, Section VIII concludes the paper and outlines directions for future work.
II Related Work
@@
Although these studies demonstrate the promise of multimodal and non-invasive stroke screening, existing methods either focus on a limited set of modalities, rely on raw identity sensitive data, or are designed primarily for clinical rather than prehospital contexts. Motivated by these limitations, the present study aims to develop a privacy aware, multimodal deep learning system that integrates facial landmarks, speech signals, and upper body movement information for early stroke screening in prehospital environments.
III Proposed Method
III-A Proposed Framework
@@
Our proposed method follows a multimodal paradigm that jointly leverages visual and auditory signals to support stroke screening in prehospital settings, as illustrated in the Fig.1. The system takes as input facial videos, upper-body videos, and speech signals collected from participants during the F.A.S.T. assessment. These data streams are processed through modality-specific branches, each tailored to the characteristics of the corresponding modality, before being integrated at the feature representation level.
@@
Figure 1: The proposed framework comprises four stages: (a) Raw Input Processing, which transforms raw visual and audio data into structured representations; (b) Modality Encoders, which extract high-level embeddings using modality-specific models; (c) Multimodal Fusion, which integrates the embeddings through a learnable fusion mechanism; and (d) Classification Head, which produces the final binary stroke prediction..
@@
In the visual branch, facial and upper-body videos are first processed using the Google MediaPipe library to extract landmark representations. For facial analysis, facial landmarks capture the geometric structure and asymmetry of critical regions such as the mouth and cheeks, which are directly associated with hallmark symptoms of stroke. In parallel, upper-body pose landmarks are employed to model the kinematics of the torso and limb movements, enabling the detection of motor abnormalities, such as impaired coordination during motor assessment tasks. Landmarks are extracted on a frame-by-frame basis and organized into temporal sequences, yielding a sequential representation of motion dynamics over time.
In the first visual sub branch, the facial landmark sequences are directly fed into an Audio Spectrogram Transformer (AST) model pretrained on ImageNet to extract spatio temporal features. Applying AST to landmark-based representations allows the model to capture long range temporal dependencies by formulating the motion sequence as a patch-based representation learning problem, analogous to image processing or time frequency analysis. In contrast, the upper-body pose landmark sequences are processed using an MLP-Mixer architecture to effectively model interactions along both spatial and temporal dimensions, which aligns well with the kinematic characteristics of body movements. Beyond its role as a lightweight feature extractor, the MLP-Mixer is particularly well suited for pose-based representations, as it explicitly disentangles token-mixing operations, corresponding to interactions among body joints, from channel-mixing operations, which capture temporal dynamics across frames. This design enables efficient learning of structured geometric relationships and motion patterns without relying on complex attention mechanisms, making it robust to variable sequence lengths and noisy landmark estimates commonly encountered in prehospital data. As a result, the model can effectively capture abnormal motion patterns, such as reduced movement amplitude or impaired coordination of the upper limbs and torso, while maintaining lower computational complexity and improved generalization. The outputs of each visual branch are subsequently mapped through projector layers to a shared embedding space with a unified dimensionality of 256.
In the auditory branch, speech signals are transformed into mel spectrograms using the short time Fourier transform applied over fixed-size sliding windows, followed by a conversion to the mel scale to emphasize frequency bands that are most informative for speech analysis. The resulting mel spectrograms are normalized in terms of sampling rate and temporal length before being passed into an AST model pretrained on ImageNet to extract high level auditory representations. These auditory features are further projected into the same embedding space as the visual modalities to facilitate multimodal integration.
@@
Once embeddings from the three modalities facial landmarks, upper-body pose, and audio are obtained, the system performs multimodal fusion using an attention-based fusion mechanism. The attention module plays a central role in coordinating and integrating information across modalities by adaptively learning the relative importance of each modality for the stroke screening task. Rather than employing simple fusion strategies such as feature concatenation or linear summation, the attention-based fusion enables the model to capture cross-modal relationships and emphasize discriminative and consistent cues while suppressing noisy or less informative signals. In practice, the reliability and expressiveness of different modalities may vary across individuals and scenarios; for instance, facial asymmetry may be more pronounced in some cases, whereas speech impairment or motor dysfunction may provide stronger indicators in others. Through learned attention weights during training, the model dynamically adjusts its fusion strategy on a per-sample basis. Consequently, the final fused representation reflects a balanced and flexible interaction among facial expressions, body kinematics, and speech characteristics over time, leading to improved robustness and predictive performance in real-world prehospital settings.
@@
Finally, the fused multimodal feature vector is fed into a classification head consisting of Layer Normalization, a GELU activation function, Dropout, and a Linear layer to perform binary classification between stroke and non-stroke conditions. The overall design of the proposed framework leverages the strengths of pretrained models while maintaining flexibility and scalability, making it suitable for practical deployment in prehospital and emergency care scenarios.
IV Dataset
IV-A Data acquisition
@@
Data collection was conducted using the rear camera of an iPhone 12 mounted on a simple tripod, positioned at a height aligned with the subject’s face while seated. To ensure consistency and minimize environmental noise, all recordings were performed in a quiet indoor environment against a neutral background (e.g., a white wall). Videos were captured in vertical orientation at a resolution of 1080p and a frame rate of 30 fps, with High Dynamic Range (HDR) disabled and mono audio recording enabled. Although not strictly required, the camera grid mode was activated to assist in centering the subject’s face within the frame.
@@
The dataset comprises recordings collected from 37 participants under a controlled prehospital setting. All participants performed a predefined set of tasks based on the F.A.S.T. assessment protocol. Detailed demographic information and data acquisition statistics are summarized in Table I.
Table I: Participant demographics and data acquisition details.
@@
Female participants 19 (51.35%)
Male participants 18 (48.65%)
Age range (18–30) 35 participants
Age range (30–40) 2 participants
Assessment protocol F.A.S.T. test
Number of task types 3
Task modalities Face, Speech, Arm movement
Healthy trials per task 3 repetitions
Simulated stroke trials per task 3 repetitions
Total trials per task 6
Total recordings per participant 18 videos
As summarized in Table II, each subject was instructed to perform three tasks corresponding to the components of the F.A.S.T. assessment. For the facial task, subjects were asked to smile as widely as possible under normal conditions. To simulate facial paralysis, they were instructed to smile using only one side of the face.
For the speech task, participants were required to repeat the sentence “Hôm nay tr`ơi đẹp quá”, which translates to “Today is such a nice day”. This sentence was selected due to its inclusion of multiple plosive phonemes, which are known to be particularly affected by dysarthria—a common speech impairment caused by disrupted neural control of oral muscles in stroke patients [2]. When simulating stroke symptoms, subjects were asked to articulate the sentence more slowly and unclearly, mimicking impaired control of the mouth and tongue muscles.
@@ IV-B Data preprocessing
Beyond facial features, body landmarks were extracted to characterize upper-body motion and skeletal dynamics, as body landmark representations have been shown to effectively capture motion patterns in video-based analysis of neurological conditions[8]. The MediaPipe Pose Landmarker (Full) model was employed, providing 33 landmarks with x, y, and z coordinates. In contrast to facial landmarks, all three axes were preserved for pose representation to retain depth information, which is particularly important for movements involving spatial displacement, such as finger to nose gestures. Collectively, these preprocessing pipelines transform each video into a temporal sequence of landmark coordinates, effectively representing motion as time-series data suitable for sequential modeling. Examples of the extracted facial and pose landmarks are illustrated in Figure2.
(a) Facial landmarks extracted using MediaPipe Face Landmarker (478 landmarks).
(b) Pose landmarks extracted using MediaPipe Pose Landmarker (33 landmarks).
Figure 2: Illustrative examples of facial and pose landmarks adapted from the Google MediaPipe documentation.
@@
Figure 3: Side-by-side comparison of mel spectrograms from (a) a healthy subject and (b) a stroke-simulated subject, illustrating differences in temporal and spectral speech patterns at 80 mel bins.
Table II: Tasks used in the F.A.S.T. assessment and their corresponding stroke-related impairments.
@@
The patient repeats the sentence “Hôm nay tr`ơi đẹp quá” (Vietnamese for “The weather is so beautiful today”) three times while looking at the camera Speech becomes slowed and pronunciation is unclear Slurred speech
@@
V Experiment
V-A Implementation Details

### v2

@@
Digital FAST: An AI-Driven Multimodal Framework for Rapid and Early Stroke Screening
NGOC-KHAI HOANG1, THI-NHU-MAI-NGUYEN3, HUY-HIEU PHAM1,2,∗ 1VinUni-Illinois Smart Health Center, VinUniversity, Hanoi 100000, Vietnam
2College of Engineering and Computer Science, VinUniversity, Hanoi 100000, Vietnam
3 Study Program Medical Informatics, Universität zu Lübeck, Germany
*Email: hieu.ph@vinuni.edu.vn
Tóm t\́abrevet n\̣ocircumflexi dung
@@ 3 Study Program Medical Informatics, Universität zu Lübeck, Germany
Keyword. Stroke detection, Multimodal Learning, Pretrained Model, Attention Fusion.
1 INTRODUCTION
Stroke refers to a group of cerebrovascular disorders caused by the formation of blood clots or hemorrhagic events that obstruct or rupture blood vessels supplying the brain. Among these conditions, acute ischemic stroke is one of the leading causes of long-term disability and mortality worldwide[18, 21], with epidemiological studies estimating that approximately one in four individuals will experience a stroke during their lifetime [10, 11, 17]. Consequently, timely and appropriate intervention is of critical importance[25]. Early detection of stroke is essential to enable prompt preventive measures and effective treatment. Notably, when individuals exhibit early signs of stroke, bystanders often lack sufficient medical knowledge to recognize these symptoms as accurately as healthcare professionals[15]. This gap highlights the urgent need for automated methods capable of detecting stroke risk in real-world settings, particularly for non-expert users. In the context of rapid digitalization, artificial intelligence (AI) has significantly impacted the healthcare domain. Deep learning models constitute the foundation of various applications[19], including automated healthcare systems, clinical decision support, and medical diagnosis, where AI has, in certain cases, achieved performance superior to that of human specialists[2, 27]. Multimodal learning approaches, which integrate information from multiple data sources, have demonstrated superior performance compared to models relying on a single modality[5, 4, 23]. However, in light of the rapid digitalization of healthcare and the growing adoption of multimodal deep learning techniques, there remains a notable gap between methodological advances and their application to practical, high-impact clinical scenarios such as prehospital stroke screening. Although prior studies have demonstrated the advantages of integrating heterogeneous data sources[1], multimodal stroke detection particularly under subject independent and prehospital settings has not yet been sufficiently explored[26, 16]. Addressing this gap requires not only effective multimodal modeling strategies but also datasets and experimental settings that reflect real-world constraints.
@@ 1 INTRODUCTION
The remainder of this paper is organized as follows. Section 2 reviews related work on stroke detection and multimodal deep learning approaches. Section 3 presents the proposed multimodal framework, including the modality specific encoders and the fusion strategy. Section 4 describes the dataset and data acquisition process in detail. Section 5 reports the experimental setup and quantitative performance results. Section 6 provides ablation studies to analyze the contribution of individual modalities and fusion strategies. Section 7 discusses the experimental findings, limitations, and practical implications of the proposed method. Finally, Section 8 concludes the paper and outlines directions for future work.
2 Related Work
@@
Further multimodal learning approaches have also been explored for rapid stroke assessment in emergency clinical settings. Yu et al.[30] proposed a multimodal deep learning framework designed to support rapid stroke diagnosis by emulating clinical screening protocols such as the Cincinnati Prehospital Stroke Scale (CPSS) and the Face Arm Speech Test (FAST). Their approach integrates facial motion analysis and speech assessment extracted from patient video recordings to automatically detect neurological deficits associated with stroke. Experiments conducted on video data collected from real patients in emergency rooms demonstrated that the proposed system achieved diagnostic performance comparable to that of emergency physicians while enabling faster automated screening.
Building upon this direction, Cai et al.[7] introduced DeepStroke, a multimodal adversarial deep learning framework for stroke screening in emergency room environments. The system jointly analyzes facial muscle coordination and speech impairments to detect subtle neurological abnormalities indicative of stroke. By leveraging multimodal feature learning and adversarial training strategies, DeepStroke effectively captures complementary information from visual and acoustic modalities. Experimental results showed that the proposed framework achieved high sensitivity for stroke detection while maintaining practical screening time suitable for emergency clinical workflows.
@@
Although these studies demonstrate the effectiveness of multimodal learning for stroke assessment, most existing systems are designed for controlled clinical environments and rely on raw facial video data, which may raise privacy concerns and limit their applicability in practical prehospital scenarios. Motivated by these limitations, the present study aims to develop a privacy-aware multimodal deep learning system that integrates facial landmarks, speech signals, and upper-body movement information for early stroke screening in prehospital environments.
3 Proposed Method
@@
We propose a multimodal deep learning framework that jointly leverages facial expressions, upper-body movements, and speech signals for automated stroke screening in prehospital settings, as illustrated in Fig. 1. The system takes as input facial videos, upper-body videos, and speech signals collected from participants during the F.A.S.T. assessment. These data streams are processed through modality-specific branches, each tailored to the characteristics of the corresponding modality, before being integrated at the feature representation level.
@@
Hình 1: The proposed framework comprises four stages: (a) Raw Input Processing, which transforms raw visual and audio data into structured representations; (b) Modality Encoders, which extract high-level embeddings using modality-specific models; (c) Multimodal Fusion, which integrates the embeddings through a learnable fusion mechanism; and (d) Classification Head, which produces the final binary stroke prediction.
3.1 Raw Input Processing
@@
As illustrated in Fig. 1(a), the system processes three types of raw inputs: facial videos, upper-body videos, and speech signals collected during the F.A.S.T. assessment task. In the visual branch, facial and upper-body videos are first processed using the Google MediaPipe library to extract landmark representations. For facial analysis, facial landmarks capture the geometric structure and asymmetry of critical regions such as the mouth and cheeks, which are directly associated with hallmark symptoms of stroke. In parallel, upper-body pose landmarks are employed to model the kinematics of the torso and limb movements, enabling the detection of motor abnormalities, such as impaired coordination during motor assessment tasks.
Landmarks are extracted on a frame-by-frame basis and organized into temporal sequences, yielding a sequential representation of motion dynamics over time. Specifically, we treat each frame as a token in a temporal sequence. For the face modality, each token is a flattened 2D landmark vector of dimension $D_{f}=478\times 2=956$; for the pose modality, each token is a flattened 3D vector of dimension $D_{p}=33\times 3=99$. A linear projection maps each per-frame token into the encoder embedding space, and positional embeddings are added to preserve temporal order. During batching, sequences are padded (or center-cropped if longer) along the temporal dimension to ensure consistent tensor shapes across samples.
For the auditory modality, speech signals are converted into mel-spectrogram representations using a short-time Fourier transform with sliding windows, followed by a transformation to the mel scale to emphasize perceptually relevant frequency bands. These spectrograms serve as the input to the speech encoder.
3.2 Modality Encoders
As shown in Fig. 1(b), each modality is processed using a dedicated encoder designed to capture modality-specific patterns.
In the first visual sub-branch, the facial landmark sequences are directly fed into an Audio Spectrogram Transformer (AST) model pretrained on ImageNet to extract spatio-temporal features. Applying AST to landmark-based representations allows the model to capture long-range temporal dependencies by formulating the motion sequence as a patch-based representation learning problem, analogous to image processing or time-frequency analysis.
In contrast, the upper-body pose landmark sequences are processed using an MLP-Mixer architecture to effectively model interactions along both spatial and temporal dimensions, which aligns well with the kinematic characteristics of body movements. Beyond its role as a lightweight feature extractor, the MLP-Mixer is particularly well suited for pose-based representations, as it explicitly disentangles token-mixing operations, corresponding to interactions among body joints, from channel-mixing operations, which capture temporal dynamics across frames.
This design enables efficient learning of structured geometric relationships and motion patterns without relying on complex attention mechanisms, making it robust to variable sequence lengths and noisy landmark estimates commonly encountered in prehospital data. As a result, the model can effectively capture abnormal motion patterns, such as reduced movement amplitude or impaired coordination of the upper limbs and torso, while maintaining lower computational complexity and improved generalization.
In the auditory branch, the mel-spectrogram representations are processed using a pretrained AST model to extract high-level auditory features that characterize speech impairments associated with stroke symptoms. The outputs of all modality encoders are subsequently mapped through projector layers to a shared embedding space with a unified dimensionality of 256. Formally, let $X_{f}$, $X_{p}$, and $X_{s}$ denote the input sequences for the face, pose, and speech modalities. The modality encoders extract feature representations as follows:
$H_{f}=\text{AST}(X_{f})$ (1)
$H_{p}=\text{MLPMixer}(X_{p})$ (2)
$H_{s}=\text{AST}(X_{s})$ (3)
The modality-specific features are then projected into a shared embedding space:
$z_{i}=\text{Projector}(H_{i}),\quad i\in\{f,p,s\}$ (4)
where $z_{f}$, $z_{p}$, and $z_{s}$ denote the projected embeddings of the face, pose, and speech modalities, respectively.
3.3 Multimodal Fusion
Once embeddings from the three modalities, namely facial landmarks, upper-body pose, and audio are obtained, the system performs multimodal fusion using a sigmoid-based gating mechanism,as illustrated in Fig. 1(c). Formally, let $z_{f}$, $z_{p}$, and $z_{s}$ denote the projected embeddings of the face, pose, and speech modalities, respectively. To adaptively modulate the contribution of each modality, a sigmoid-based gating mechanism is applied to each embedding.
$z^{\prime}_{f}=z_{f}\odot\sigma(z_{f})+z_{f}$ (5)
$z^{\prime}_{p}=z_{p}\odot\sigma(z_{p})+z_{p}$ (6)
$z^{\prime}_{s}=z_{s}\odot\sigma(z_{s})+z_{s}$ (7)
where $\sigma(\cdot)$ denotes the sigmoid activation function and $\odot$ represents element-wise multiplication. This gating operation adaptively enhances informative modality features while suppressing less relevant signals.
The final multimodal representation is obtained by aggregating the gated embeddings:
$z^{\prime}=z^{\prime}_{f}+z^{\prime}_{p}+z^{\prime}_{s}$ (8)
where $z^{\prime}$ denotes the fused multimodal feature vector used for downstream classification.
The attention module plays a central role in coordinating and integrating information across modalities by adaptively learning the relative importance of each modality for the stroke screening task. Rather than employing simple fusion strategies such as feature concatenation or linear summation, the attention-based fusion enables the model to capture cross-modal relationships and emphasize discriminative and consistent cues while suppressing noisy or less informative signals.
@@ 3.3 Multimodal Fusion
In practice, the reliability and expressiveness of different modalities may vary across individuals and scenarios; for instance, facial asymmetry may be more pronounced in some cases, whereas speech impairment or motor dysfunction may provide stronger indicators in others. Through learned attention weights during training, the model dynamically adjusts its fusion strategy on a per-sample basis. Consequently, the final fused representation reflects a balanced and flexible interaction among facial expressions, body kinematics, and speech characteristics over time, leading to improved robustness and predictive performance in real-world prehospital settings.
3.4 Classification Head
@@
Finally, as shown in Fig. 1(d), the fused multimodal feature vector is fed into a classification head consisting of Layer Normalization, a GELU activation function, Dropout, and a Linear layer to perform binary classification between stroke and non-stroke conditions.
Given the fused representation $z^{\prime}$, the classification head first produces an intermediate feature representation through normalization and nonlinear transformation:
$h=\text{GELU}(\text{LayerNorm}(z^{\prime}))$ (9)
This representation is then passed through a linear classifier to compute the prediction score
$y=\text{Linear(Dropout(}(h))$ (10)
and a sigmoid activation function is applied to obtain the final probability of the stroke class
$\hat{y}=\text{Sigmoid}(y)$ (11)
The overall design of the proposed framework leverages the strengths of pretrained models while maintaining flexibility and scalability, making it suitable for practical deployment in prehospital and emergency care scenarios.
4 Dataset
4.1 Data acquisition
@@
Data collection was conducted using the rear camera of an iPhone 12 mounted on a simple tripod, positioned at a height aligned with the subject’s face while seated. An overview of the recording setup and task-specific framing is illustrated in Fig. 2. To ensure consistency and minimize environmental noise, all recordings were performed in a quiet indoor environment against a neutral background. Videos were captured in vertical orientation at a resolution of 1080p and a frame rate of 30 fps, with High Dynamic Range (HDR) disabled and mono audio recording enabled. Although not strictly required, the camera grid mode was activated to assist in centering the subject’s face within the frame.
Hình 2: (a) Standardized data acquisition setup using a rear iPhone 12 camera mounted on a tripod at face height while the subject was seated. (b) Recording tasks including speech, upper-body pose, and facial expression assessment.
@@
The dataset comprises recordings collected from 37 participants under a controlled prehospital setting. We define one multimodal sample (trial) as a triplet of recordings (Face, Speech, and Pose) captured for the same repetition. Each participant performed 6 repetitions (3 healthy + 3 simulated stroke), resulting in 6 multimodal samples per participant and 222 multimodal samples in total. Each multimodal sample contains three modality-specific videos, yielding 666 raw video clips $(222\times 3)$ before quality filtering. All participants performed a predefined set of tasks based on the F.A.S.T. assessment protocol. Detailed demographic information and data acquisition statistics are summarized in Table 1.
[…]
@@
Female / Male 19 (51.35%) / 18 (48.65%)
Age (18–30) / (30–40) 35 / 2
Protocol and modalities
Assessment protocol F.A.S.T.-inspired tasks
Task types 3 (Face, Speech, Arm movement)
Modalities per sample 3 (Face, Speech, Pose)
Recording counts
Repetitions per condition (per participant) 3 healthy + 3 simulated stroke
Total repetitions (per participant) 6
Multimodal samples (per participant) 6
Total multimodal samples 222
Raw video clips (per participant) 18 (6 samples $\times$ 3 modalities)
Total raw video clips 666 (222 samples $\times$ 3 modalities)
As summarized in Table 2, each subject was instructed to perform three tasks corresponding to the components of the F.A.S.T. assessment. For the facial task, subjects were asked to smile as widely as possible under normal conditions. To simulate facial paralysis, they were instructed to smile using only one side of the face.
For the speech task, participants were required to repeat the sentence “Hôm nay trời đẹp quá”, which translates to “Today is such a nice day”. This sentence was selected due to its inclusion of multiple plosive phonemes, which are known to be particularly affected by dysarthria - a common speech impairment caused by disrupted neural control of oral muscles in stroke patients [3]. When simulating stroke symptoms, subjects were asked to articulate the sentence more slowly and unclearly, mimicking impaired control of the mouth and tongue muscles.
@@ 4.2 Data preprocessing
Beyond facial features, body landmarks were extracted to characterize upper-body motion and skeletal dynamics, as body landmark representations have been shown to effectively capture motion patterns in video-based analysis of neurological conditions[12]. The MediaPipe Pose Landmarker[14] model was employed, providing 33 landmarks with x, y, and z coordinates. In contrast to facial landmarks, all three axes were preserved for pose representation to retain depth information, which is particularly important for movements involving spatial displacement, such as finger to nose gestures. Collectively, these preprocessing pipelines transform each video into a temporal sequence of landmark coordinates, effectively representing motion as time-series data suitable for sequential modeling. Examples of the extracted facial and pose landmarks are illustrated in Figure 3.
(a) Facial landmarks extracted using MediaPipe Face Landmarker[13] (478 landmarks).
(b) Pose landmarks extracted using MediaPipe Pose Landmarker[14] (33 landmarks).
Hình 3: Illustrative examples of facial and pose landmarks adapted from the Google MediaPipe documentation.
@@
Hình 4: Side-by-side comparison of mel spectrograms from (a) a healthy subject and (b) a stroke-simulated subject, illustrating differences in temporal and spectral speech patterns at 80 mel bins.
[…]
@@
The patient repeats the sentence “Hôm nay trời đẹp quá” (Vietnamese for “The weather is so beautiful today”) three times while looking at the camera Speech becomes slowed and pronunciation is unclear Slurred speech
@@
5 Experiments
5.1 Implementation Details